# Inherent and emergent liability issues in LLM-based agentic systems: a principal-agent perspective

### v1

@@ Abstract
Agentic systems powered by large language models (LLMs) are becoming progressively more complex and capable. Their increasing agency and expanding deployment settings attract growing attention over effective governance policies, monitoring and control protocols. Based on emerging landscapes of the agentic market, we analyze the potential liability issues stemming from delegated use of LLM agents and their extended systems from a principal-agent perspective. Our analysis complements existing risk-based studies on artificial agency and covers the spectrum of important aspects of the principal-agent relationship and their potential consequences at deployment. Furthermore, we motivate method developments for technical governance along the directions of interpretability and behavior evaluations, reward and conflict management, and the mitigation of misalignment and misconduct through principled engineering of detection and fail-safe mechanisms. By illustrating the outstanding issues in AI liability for LLM-based agentic systems, we aim to inform the system design, auditing and monitoring approaches to enhancing transparency and accountability.
@@ 1 Introduction
AI agents are computer software systems capable of creating context-specific plans in non-deterministic environments (Chan et al., 2023; Krishnan, 2025). AI agents based on LLMs (aka. LLM agents, see Appendix A) exist on a spectrum of autonomy, ranging from simple tool-calling agents to generalist agents capable of planning, sourcing, critiquing, and executing their own workflow (Li, 2025). They primarily adopt an architecture with explicitly defined functioning components [^1] Also called cognitive architecture (Kotseruba & Tsotsos, 2020) at times, but the architecture alone doesn’t guarantee cognition or sense of agency. (Sumers et al., 2023). LLM-based multiagent systems (MASs) allow agents to interact, collaborate, or compete within shared environments (Fig. 1a). They are designed by combining agents with specialized roles through LLM role-playing (Shanahan et al., 2023; Chen et al., 2024a) or through integration on a software platform. Each agent handles specific subtasks based on its expertise and allocated resources (tools, data, compute, etc). MAS can be tailored to a wide range of scales and domains, from few-agent systems that simulate team decision-making in medicine (Tang et al., 2024; Kim et al., 2024) and finance (Xiao et al., 2025), to many-agent systems that mimic the population-level socioeconomic dynamics (Park et al., 2024a; Piao et al., 2025). A plausible LLM-based MAS (Fig. 1a) can possess multiple teams of interacting agents coordinated by an orchestrator (Wang et al., 2025) and behaviorally regulated by other platform agents or through a set of engineered constraints (e.g. norms) (Criado et al., 2011; Hadfield-Menell et al., 2019). While the flexibility of LLM-based MAS provides adaptability to various applications and additional affordances but also emergent risks not present in single agents (Hammond et al., 2025; Pan et al., 2025).
@@
Figure 1: (a) A plausible LLM-based MAS deployed on an agent platform, where delegation goes from the principal to an orchestrator (agent) to different functioning agent teams. The platform also contains supporting agents. Black circles indicate teams. (b) Landscape of the LLM agentic market. LLM agents are supplied through agent-native services or agentic SaaS (agents on top of existing software apps). Agent suppliers are supported by model suppliers. Colors in (a)-(b) distinguish between agents of different types or suppliers.
@@
Understanding the current landscape of the agentic market [^2] https://aiagentslist.com/ai-agents-map (Fig. 1b and Appendices A-B) is essential for analyzing liability. The present work uses the following terms to refer to its key components: Software platforms offering agent-native services include generalist agents that focuses on general autonomous use of computers, specialist agents which target labor intensive sectors and provide verticalized services for domain-specific workflow automation (Bousetouane, 2025), and character-infused, “hireable” agentic employees. Specialist agents are also offered directly by established software platforms as agentic SaaS to streamline their existing app services. Both agent-native services and agentic SaaS largely source models externally from model suppliers or derive their own models from open-source projects. Separately, prototypes of integration platforms and integration protocols facilitating the interaction of agents from different frameworks with third-party resources during deployment are also appearing.
@@
Although the potential liability issues in the rapidly expanding market are prevalent, the governance of AI agents is still a nascent topic. Existing efforts are built along two streams: one focuses on establishing and refining the taxonomies of risks and harms using empirical evidence (Chan et al., 2023; 2024; He et al., 2024; Hammond et al., 2025) or differentiating the governance of agentic systems from traditional machine learning (ML) models (Cohen et al., 2024; Kolt, 2025); another focuses on understanding the interactions between humans and AI agents to build constructive principles (Zheng et al., 2023). Research in both streams used principal-agent theory (PAT) (Eisenhardt, 1989; Laffont & Martimort, 2002) as a starting point, but lacked systematic examination of how existing legal liability frameworks can effectively address various principal-agent relationships in AI systems. Moreover, AI systems are typically embedded in a sociotechincal system (Weidinger et al., 2023) such that the risks are materialized only through interacting with the environment (see example and analysis in Appendix C).
@@
Despite the long history of PAT-based legal frameworks (Munday, 2022), their use in AI systems is still in early days. This work presents an initial attempt in extending principal-agent analysis to current LLM-based agentic systems. Our major contributions include: (i) extending the previous work on PAT for AI governance to accommodate the characteristic behaviors of LLM agents; (ii) discussion of the potential liability issues in light of the emerging market for LLM-based agentic systems; and (iii) outlining of the important directions of policy-driven method developments that support the technical governance of LLM agents.
@@ 2 Related works
Legal scholars have treated AI as products and the harm they caused under product liability (Abraham & Rabin, 2019; Buiten, 2024; Sharkey, 2024). In the US, product liability falls under: (i) design defect, (ii) manufacturing defect, and (iii) a warning defect (inadequate instructions) (Abraham, 2012). Anyone harmed by a product can bring a claim. Product harm carries strict liability — meaning courts do not consider the care level — and affects anyone along the supply chain.
Service-caused harm usually trigger a negligence inquiry: was “reasonable care” used? If not, the service provider is liable (primary liability) and their principal can also be liable under multiple legal theories. Principals face two primary liabilities: “negligent hiring” for hiring an agent without a reasonable due diligence on its modus operandi; and “negligent supervision” for failing to reasonably monitor or control its agents. Principals face secondary liability such as vicarious liability (Sykes, 1984; Diamantis, 2023) because of their relationship (when its agents acted within the boundaries of employment). A service relationship can also fall under a contractor relationship, which does not trigger vicarious liability. Courts look at the contract and control exercised to categorize the relationship. Below, we assume the principal exercise enough control to trigger vicarious liability.
Software exhibits both “product” (movable goods) and “service” (akin to professional offering) characteristics (Gemignani, 1980; Popp, 2011). As AI agents become more autonomous, they move closer to services, their actions are more accountable due to increasing agenticness (Chopra & White, 2011; Chan et al., 2023). A negligence rule with potential vicarious liability may be more suited (Turner, 2018) for those relationships. In current legal use, this framework straddles product liability falling on the manufacturer and vicarious liability involving principals. Besides the service-product divide, law and economics (L&E) has also advocated other approaches, including risk-based liability (Geistfeld et al., 2022), fault-based liability (Buiten et al., 2023), explanation-based liability (Padovan et al., 2023), etc.
@@
Prior works that invoked the principal-agent framework in AI/ML primarily focused on the decision-making aspects in human-AI collaboration and AI safety. Lubars & Tan (2019) discussed the relation between task delegation and the principal’s preference. Hadfield-Menell & Hadfield (2019) mapped AI alignment onto the principal-agent problem and discussed the alignment issues in the incomplete contracts theory. Athey et al. (2020) considered different scenarios in allocating decision authority when the human principal and the AI agent have different capabilities. Critch & Krueger (2020) and Hendrycks (2023) considered the potential dangers of complete task delegation to agentic systems. Besides, principal-agent problems have also been considered in game-theoretic machine learning (Gan et al., 2024) and in reinforcement learning settings with two interacting agents (Ivanov et al., 2024).
@@
Agentic systems have long been proposed as a canonical approach for software design (Jennings, 2001; Zambonelli et al., 2003). Their core advantages are the scalability, flexibility, and the ability to perform complex tasks through task decomposition. LLM agents (Li, 2025) can be configured by text instruction and can interact with external resources to achieve enhanced capabilities than LLMs in reasoning, tool use, memory, planning, and can be personalized according to the role-playing capability of the base model (Shanahan et al., 2023; Chen et al., 2024a). LLM-based MASs can use verbal communication protocols to facilitate collaboration and engage in debates (Tran et al., 2025), and the protocol topology is a key for their efficient scaling and behavioral control (Qian et al., 2025).
3 Principal agency and liability
PAT examines the relationship where one party, the principal, delegates authority to another, the agent, creating three fundamental challenges (Eisenhardt, 1989; Laffont & Martimort, 2002):
  3. Misaligned interest (aka. conflict of interest) can manifest between principals and agents in four ways (Jensen & Meckling, 1976): principal-agent collusion against third parties, principal-third party collusion against agents, agent-third party collusion against principals, or agents simply pursuing self-interest independently. Misaligned interests lead to agency cost, which is associated with information sharing, monitoring of the agent, etc (Fama & Jensen, 1983).
Each of these problems has legal and economic solutions. Legal solutions include fiduciary duties (duty of care, duty of loyalty, etc.). Those duties provide principals with a legal recourse against agents who breached those duties. However, legal enforcement is probabilistic and slow (Levmore, 1990), so principals often prefer to use economic mechanisms. Economic solutions depend on the problem. Principals can address some hidden information problems by creating separating equilibria that force agents to reveal information about themselves, that is, “signals”; signals include credentials, warranties, or performance histories that distinguish high-quality from low-quality agents (Spence, 1973). Principals mitigate a hidden action problem through monitoring (i.e. direct observation of agent behavior) and then linking compensation to observable effort, or bonding arrangements in which the principals tie remuneration to outcomes (bonuses), thereby aligning financial incentives (Holmström, 1979). A principal can address conflict of interest by realigning incentives through carefully designed contracts (deferred compensation), creating organizational structures that promote incentive realignment (profit sharing), or leveraging reputation mechanisms. A principal has incentives to monitor their agents. Primary and secondary liability exposures provide additional incentives to ensure that it oversees that its agents complete their tasks using reasonable care.
@@ 4 Inherent liability issues in single agents
Contemporary approaches to the governance of AI agents (Kampik et al., 2022; Chan et al., 2023; 2024; Kolt, 2025) resort to PAT, where the principal, the human or company, delegates a task or goal to the AI agent, based on a mutual agreement. Yet LLM agents still cannot satisfy all criteria of a normal agent (Perrier & Bennett, 2025) in PAT, therefore, the agency gap can create additional liability issues or liability gaps (Turner, 2018).
@@ 4.1 Artificial agency
PAT requires a clarification of the agency relationship, which remains a hotly debated interdisciplinary topic for LLM-based systems (Shavit et al., 2023; Dai, 2024; Barandiaran & Almendros, 2024; Dung, 2024; Perrier & Bennett, 2025; Mattingly & Cibralic, 2025; Butlin, 2025; Das, 2025). AI researchers often take an operational view of artificial agency, such that it is possible to quantify and compare the mental state characteristics of and between AI systems through external interrogation (Baird & Maruping, 2021; Chan et al., 2023; Miehling et al., 2025). Representative caveats of artificial agency for LLMs and LLM agents include:
@@
These caveats finds similar manifestations in behavioral economics (Thaler, 2016).
@@
Because an agency relationship requires both principals and agents agree to the relationship, current LLM agents cannot yet form an authentic relationship of such because of their flawed agency (Barandiaran & Almendros, 2024; Perrier & Bennett, 2025). Voluntary relationships usually indicates that both parties benefit. However, because rationality is generally not a built-in goal in developing LLM agents (Macmillan-Scott & Musolesi, 2024), their lack of consent (from agents) means that agent suppliers may face liability or risks they have not considered. That is why scholars have consistently argued that agent failures should be treated as product liability, and the allocation should reflect that (Abraham & Rabin, 2019; Buiten, 2024; Sharkey, 2024). In such a situation, the Coase Theorem suggests that the allocation of liability and risks is best left to contracts but before the parties can do that, legislators and courts must clarify rights and liabilities (Coase, 1960).
@@ 4.2 Task specification and delegation
Principals delegate tasks to their agent usually either because the principals lack the resources or the expertise to deal with the tasks (Castelfranchi & Falcone, 1998). For LLM agents, the principal provides as input to the LLM a task specification that contains instructions on the nature and procedure of task execution, available resources, potential ways to overcome hurdles, the principal’s preferences, etc. Such a procedure is subject to
  1. Task underspecification is often present because it is impossibly costly to fully anticipate all possible scenarios to put into the specification. Moreover, the same agent may be appropriate for one task but not another; therefore, task specification — the equivalent of job design in organizational theory (Oldham & Fried, 2016) — should balance the tradeoff between agent capability and task complexity (Hadfield-Menell & Hadfield, 2019). This incompleteness leaves the door open to undesirable outcomes such as negative side effects.
  2. Risky delegation The principal tends to forgo the delegation of very high-stakes tasks because of their severe adverse consequences and the unreliability of agent behaviors (Lubars & Tan, 2019). The risk of delegation can be reduced by the amount of repeated feedback the principal provides to ensure alignment (Jiang et al., 2024).
@@
When selecting the agent (and the tasks) to delegate, principals are expected to carry out their due diligence. If they fail to take reasonable precautions when selecting agents, principals retain liability for negligent hiring (Camacho, 1993) and negligent retention. At the best of time, delegation of tasks raises information concerns because the principal may not have authorized the agent to share information. In the LLM context, the sharing of information raises concerns of copyright and trademark (e.g. duplication of protected documents), trade secret (sharing information outside the system), privacy (e.g. transferring data governed by the General Data Protection Regulation or California Consumer Privacy Act), etc. That is why many tasks are not delegable. A human agent may be able to dissociate the difference between delegable and non-delegable tasks based on the sensitivity of the information, while an AI agent may not without clear instructions (Hadfield-Menell & Hadfield, 2019).
Figure 2: Examples of interaction patterns between the principals and single agents or multiagent systems (MASs). Each MAS in the coordination or collaboration (collab.)/competition (compet.) pattern is enclosed within a box.

### v2

@@ Abstract
Agentic systems powered by large language models (LLMs) are becoming progressively more complex and capable. Their increasing agency and expanding deployment settings attract growing attention to effective governance policies, monitoring, and control protocols. Based on the emerging landscape of the agentic market, we analyze potential liability issues arising from the delegated use of LLM agents and their extended systems through a principal-agent perspective. Our analysis complements existing risk-based studies on artificial agency and covers the spectrum of important aspects of the principal-agent relationship and their potential consequences at deployment. Furthermore, we motivate method developments for technical governance along the directions of interpretability and behavior evaluations, reward and conflict management, and the mitigation of misalignment and misconduct through principled engineering of detection and fail-safe mechanisms. By illustrating the outstanding issues in AI liability for LLM-based agentic systems, we aim to inform the system design, auditing, and tracing to enhance transparency and liability attribution.
Inherent and emergent liability issues in LLM-based agentic systems:
a principal-agent perspective
Garry A. Gabison1, 2, *, R. Patrick Xian3, 4, *, 1Queen Mary University of London, 2UC Berkeley, 3Certivize AI, 4UC San Francisco *Equal contribution 🖂: g.gabison @ qmul.ac.uk, xrpatrick @ gmail.com
@@ 1 Introduction
AI agents are computer software systems capable of creating context-specific plans in non-deterministic environments (Chan et al., 2023; Krishnan, 2025). AI agents based on LLMs (aka. LLM agents, see Appendix A) exist on a spectrum of autonomy, ranging from simple tool-calling agents to generalist agents capable of planning, sourcing, critiquing, and executing their own workflow (Li, 2025). They primarily adopt an architecture with explicitly defined functioning components [^1] Also called cognitive architecture (Kotseruba and Tsotsos, 2020) at times, but the architecture alone doesn’t guarantee cognition or sense of agency. (Sumers et al., 2023). LLM-based multiagent systems (MASs) allow agents to interact, collaborate, or compete within shared environments (Fig. 2a). They are designed by combining agents with specialized roles through LLM role-playing (Shanahan et al., 2023; Chen et al., 2024a) or through integration on a software platform.
@@
Figure 1: A plausible LLM-based MAS deployed on an agent platform, where delegation goes from the principal to an orchestrator (agent) to different functioning agent teams (circles). The platform also contains supporting agents for safety, security, and compliance. Colors distinguish between agents of different types.
Each agent handles specific subtasks based on its expertise and allocated resources (tools, data, compute, etc). MASs can be tailored to a wide range of scales and domains, from few-agent systems that simulate team decision-making in medicine (Tang et al., 2024; Kim et al., 2024) and finance (Xiao et al., 2025), to many-agent systems that mimic the population-level socioeconomic dynamics (Park et al., 2024a; Piao et al., 2025). A plausible LLM-based MAS (Fig. 2a) can possess multiple teams of interacting agents coordinated by an orchestrator (Wang et al., 2025) and behaviorally regulated by other platform agents or through a set of engineered constraints (e.g. norms) (Criado et al., 2011; Hadfield-Menell et al., 2019). The components mentioned here are further defined and explained in Appendix A. While the flexibility of LLM-based MASs allows adaptation to various applications and affordances, it also introduces emergent risks not present in single agents (Hammond et al., 2025; Pan et al., 2025).
Figure 2: Landscape of the LLM agentic market. LLM agents are provided through agent-native services or agentic SaaS (agents on top of existing software apps). Agent providers are supported by model providers. Colors distinguish agents of different types.
@@
Understanding the current landscape of the agentic market [^2] https://aiagentslist.com/ai-agents-map (Fig. 2b and Appendices A-B) is essential for analyzing liability. The present work uses the following terms to refer to its key components: Software platforms offering agent-native services include generalist agents that focuses on general autonomous use of computers, specialist agents which target labor intensive sectors and provide verticalized services for domain-specific workflow automation (Bousetouane, 2025), and character-infused, “hireable” agentic employees. Specialist agents are also offered directly by established software platforms as agentic SaaS to streamline their existing app services. Both agent-native services and agentic SaaS largely source models externally from model providers or derive their own models from open-source projects. Separately, prototypes of integration platforms and integration protocols facilitating the interaction of agents from different frameworks with third-party resources during deployment are also appearing.
@@
Examples of realistic agentic systems exist across many application settings (see Table 1). Because the governance of AI agents remains a nascent topic, potential liability issues in the rapidly expanding agentic market are prevalent but not sufficiently analyzed. Existing efforts are built along two streams: one focuses on establishing and refining the taxonomies of risks and harms using empirical evidence (Chan et al., 2023, 2024; He et al., 2024; Hammond et al., 2025) or differentiating the governance of agentic systems from traditional machine learning (ML) models (Cohen et al., 2024; Kolt, 2025); another focuses on understanding the interactions between humans and AI agents to build constructive principles (Zheng et al., 2023). Research in both streams used principal-agent theory (PAT) (Eisenhardt, 1989; Laffont and Martimort, 2002) as a starting point (see Appendix C), but lacked a systematic examination of how existing legal frameworks for liability can effectively address various principal-agent relationships in AI systems. Moreover, AI systems are typically embedded in a sociotechnical system (Weidinger et al., 2023) such that the verdict on liability issues are finalized only through understanding the agent-environment interactions (see example and analysis in Appendix E). LLM agents are still yet to be mass-deployed, so the liability issues we raise here are based on existing behavioral traits studied in technical research and hypotheticals (see Appendix D.3) extrapolated from them.
@@
Despite the long history of PAT-based legal frameworks (Munday, 2022), their use in AI systems is still in its early days. This work presents an initial attempt to bring principal-agent analysis to current LLM-based agentic systems. Our major contributions include: (i) extending the previous work on PAT for AI governance to accommodate the characteristic behaviors of LLM agents; (ii) discussion of the potential liability issues in light of the emerging market for LLM-based agentic systems; and (iii) outlining of the important directions of policy-driven method developments that support the technical governance of LLM agents. Without loss of generality, our analysis in the present work focuses on US jurisdictions using the Restatement of Torts and the Restatement of Law (see Appendix D) as the starting point to guide interpretation because tort and agency laws are not harmonized across the US. The framework put forward in this article can be applied to other jurisdictions by contextualizing the relevant agent behavior within the corresponding legal frameworks.
@@ 2 Related works
Liability refers to the legal responsibility for one’s actions. Legal scholars have treated AI as products and the harm they caused under product liability (Abraham and Rabin, 2019; Buiten, 2024; Sharkey, 2024). In the US, product liability falls under: (i) design defect, (ii) manufacturing defect, and (iii) a warning defect (inadequate instructions) (Abraham, 2012). Anyone harmed by a product can make a claim against any link in the supply chain. Usually, product liability is treated under a strict liability rule, courts do not consider the care level that a manufacturer put in place to avoid an accident (see Appendix D for legal definitions). Design defect can be treated under a strict liability rule or an inquiry that also resembles a negligence inquiry, courts consider the care level by considering alternative existing designs.
Service-caused harm usually triggers a negligence inquiry: was “reasonable care” used? If not, the service provider is liable (primary liability) and their principal can also be liable under multiple legal theories. Principals face two primary liabilities: “negligent hiring” for hiring an agent without a reasonable due diligence on its modus operandi; and “negligent supervision” for failing to reasonably monitor or control its agents. Principals face secondary liability, such as vicarious liability (Sykes, 1984; Diamantis, 2023), because of their relationship (when agents acted within the boundaries of employment). A service relationship can also fall under a contractor relationship, which does not trigger vicarious liability. Courts look at the contract and control exercised to categorize the relationship. Below, we assume the principal exercises enough control to trigger vicarious liability.
Software exhibits both “product” (movable goods) and “service” (akin to professional offering) characteristics (Gemignani, 1980; Popp, 2011). As AI agents become more autonomous, they move closer to services, their actions are more accountable due to increasing agenticness (Chopra and White, 2011; Chan et al., 2023). A negligence rule with potential vicarious liability may be more suited (Turner, 2018) for those relationships. In current legal use, this framework straddles product liability falling on the manufacturer and vicarious liability involving principals. Besides the service-product divide, law and economics (L&E) has also advocated for other approaches, including risk-based liability (Geistfeld et al., 2022), fault-based liability (Buiten et al., 2023), explanation-based liability (Padovan et al., 2023), etc.
@@
Prior works that invoked the principal-agent framework in AI/ML primarily focused on the decision-making aspects in human-AI collaboration and AI safety. The agents here act as representatives of the principal, which differ from the agents in traditional reinforcement learning (Diaz et al., 2024). Lubars and Tan (2019) discussed the relation between task delegation and the principal’s preference. Hadfield-Menell and Hadfield (2019) mapped AI alignment onto the principal-agent problem and discussed the alignment issues in the incomplete contracts theory. Athey et al. (2020) considered different scenarios in allocating decision authority when the human principal and the AI agent have different capabilities. Critch and Krueger (2020) and Hendrycks (2023) considered the potential dangers of complete task delegation to agentic systems. Besides, principal-agent problems have also been considered in game-theoretic machine learning (Gan et al., 2024) and in reinforcement learning settings with two interacting agents (Ivanov et al., 2024).
@@
Agentic systems have long been proposed as a canonical approach for software design (Jennings, 2001; Zambonelli et al., 2003). This paradigm has experienced further rise in the LLM era (Wang et al., 2024c) because of the role-playing capability (Shanahan et al., 2023; Chen et al., 2024a) of the these models. Their core advantages are scalability, flexibility, and the ability to perform complex tasks through task decomposition. LLM agents (Li, 2025) are configured by text instructions and they can interact with external resources to achieve enhanced capabilities than LLMs in reasoning, tool use, memory, planning, and personalization. LLM-based MASs can use verbal communication protocols to facilitate collaboration and engage in debates (Tran et al., 2025), and the protocol topology is a key for their efficient scaling and behavioral control (Qian et al., 2025).
Principal Agentic system Delegated tasks
Radiologist Medical imaging AI agent(s) Produce a preliminary interpretation of a CT/MRI scan for a patient, suggest additional tests or treatments
Frontend engineer Website design AI agent(s) Produce the code for a website from a design specification of each webpage and external media resources
Traveler Travel booking AI agent(s) Plan a trip to a series of destinations including the selection of lodgings and transportation vehicles
Shopper Online shopping AI agent(s) Seek and aggregate the best online deals that match items from a purchase list within a purchase budget
Insurance policyholder Insurance claim AI agent(s) Combine different sources of information (hospital bills, accident reports, personal emails, etc) to draft an insurance claim following a specific format restriction
Table 1: Examples of AI agent use cases cast in the principal-agent framework. Each agentic system can be implemented using a single agent or multiple agents.
@@ 3 Inherent liability issues in single agents
Contemporary approaches to the governance of AI agents (Kampik et al., 2022; Chan et al., 2023, 2024; Kolt, 2025) resort to PAT, where the principal, the human or company, delegates a task or goal to the AI agent, based on a mutual agreement. Yet LLM agents cannot satisfy all criteria of a normal agent (Perrier and Bennett, 2025) in PAT, creating an agency gap that can lead to an excess of unpredictable actions (John et al., 2025). Inherent liability issues in agentic systems arise from the dependence structure between the principal and the agent in task delegation as well as the agency gap between LLM agents and normal human agents. We discuss these issues from the perspective of each key component of PAT (see Section C) in the single-principal and single-agent setting.
@@ 3.1 Artificial agency
PAT requires a clarification of the agency relationship, which remains a hotly debated interdisciplinary topic for LLM-based systems (Shavit et al., 2023; Dai, 2024; Barandiaran and Almendros, 2024; Dung, 2024; Perrier and Bennett, 2025; Mattingly and Cibralic, 2025; Butlin, 2025; Das, 2025). AI researchers often take an operational view of artificial agency, such that it is possible to quantify and compare the mental state characteristics of and between AI systems through external interrogation (Baird and Maruping, 2021; Chan et al., 2023; Miehling et al., 2025). Representative caveats of artificial agency for LLMs and LLM agents include:
@@
An agency relationship requires both principals’ and agents’ agreement. Current LLM agents cannot yet form an authentic relationship of such because of their flawed agency (Barandiaran and Almendros, 2024; Perrier and Bennett, 2025). Voluntary relationships usually indicate that both parties benefit. However, because rationality is generally not a built-in goal in developing LLM agents (Macmillan-Scott and Musolesi, 2024), their lack of consent (from agents) means that agent providers may face liability or risks they have not considered. That is why many have consistently argued that AI failures should be treated as product liability and because of its opacity, those failures should be adjudicated under a strict liability rule (Abraham and Rabin, 2019; Buiten, 2024; Sharkey, 2024). AI providers would be left on the hook. This approach would align with the least-cost avoidance theory of assigning liability based on who can avoid the same accident at the lowest costs (Calabresi and Melamed, 1972) because users may not be the best place to assess whether the AI agents have any defects that could cause an accident. Even if they can, they may not be able to modify the functioning of the AI agents to avoid an accident.
The problem is that AI providers may not want to be held liable because they do not want to be exposed to unquantifiable risks as they cannot anticipate how the AI users will deploy their AI. In such a situation, the Coase Theorem suggests that the allocation of liability and risks is best left to contracts but before the parties can do that, legislators and courts must clarify rights and liabilities (Coase, 1960). Both the AI providers and users need to use a contract to apportion possible liability or compensation mechanisms.
@@ 3.2 Task specification and delegation
A principal delegates tasks to their agents usually either because the principal lacks the resources or the expertise to deal with the tasks (Castelfranchi and Falcone, 1998; Mitchell, 2021). For LLM agents, the principal provides as input to the LLM a task specification that contains instructions on the nature and procedure of task execution, available resources, potential ways to overcome hurdles, the principal’s preferences, etc. Such a procedure is subject to the following issues:
  1. Task underspecification is often present because it is impossibly costly to fully anticipate all possible scenarios to put into the specification. Moreover, the same agent may be appropriate for one task but not another; therefore, task specification – the equivalent of job design in organizational theory (Oldham and Fried, 2016) – should balance the tradeoff between agent capability and task complexity (Hadfield-Menell and Hadfield, 2019). This incompleteness leaves the door open to undesirable outcomes such as negative side effects.
  2. Risky delegation The principal tends to forgo the delegation of very high-stakes tasks because of their severe adverse consequences and the unreliability of agent behaviors (Lubars and Tan, 2019). The risk of delegation can be reduced by the amount of repeated feedback the principal provides to ensure alignment (Jiang et al., 2024).
@@
When selecting the agent (and the tasks) to delegate, principals are expected to carry out their due diligence. If they fail to take reasonable precautions when selecting agents, principals can face liability for negligently selecting (Camacho, 1993). Negligent selection (or negligent hiring) occurs when a principal fails to exercise reasonable care when hiring an agent and the failure caused a third party to suffer harm. Negligent selection applies beyond delegating risky tasks. When the human principal selects the original AI agent, the original selection can trigger a negligent selection. Once the AI agent starts selecting other AI agents to carry out subtasks, that selection is more likely to fall under a theory of negligent supervision (see Section 3.3).
Delegation of tasks also raises information concerns because the principal may not have authorized the agent to share information (Baird and Maruping, 2021). Granting LLM agents access to critical information raises concerns of copyright and trademark (e.g. duplication of protected documents), trade secret (sharing information outside the system), privacy (e.g. transferring data governed by the General Data Protection Regulation or California Consumer Privacy Act), etc. That is why many tasks are not delegable (Mitchell, 2021; Mitchell et al., 2025). A human agent may be able to distinguish between delegable and non-delegable tasks based on the sensitivity of the information, while an AI agent may not without clear instructions (Hadfield-Menell and Hadfield, 2019).
Figure 3: Examples of interaction patterns between the principals and single agents or multiagent systems (MASs). Each MAS in the coordination or collaboration (collab.)/competition (compet.) pattern is enclosed within a box. Distinct principals are colored differently.